OpenStudy (anonymous):

Solve, using the substitution method. j + k = 3 j – k = 7 A. The solution is (8, 1) B. The solution is (5, –2) C. There is no solution. D. There are an infinite number of solutions.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

b

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thxs so much

OpenStudy (cwrw238):

from the second equation: j = k + 7 now substitute this for j in the first equation to find value of k

OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 4x – 6 8x – 2y = 14 A. (14, 12) B. (12, 14) C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (cwrw238):

plz don't just give the answer - read the website policy - you should guide the person to the solution

OpenStudy (anonymous):

D

OpenStudy (anonymous):

substitute y into the second equation and you will get 12 = 14 which is not true....so there are no possible solution

OpenStudy (anonymous):

Solve, using the substitution method. 3x + 2y = 11 x – 2 = – 4y

OpenStudy (anonymous):

(4, 0.5)

OpenStudy (cwrw238):

am I talking to the wall ??!!

OpenStudy (anonymous):

shhh your being rude im only 9